OpenStudy (anonymous):

And i need help on this problem. Its asking to solve the linear system by using substitution. -x+2y=-6 8x+y=

OpenStudy (anonymous):

Was the earlier answer, x=-4 and y=1 incorrect?

OpenStudy (anonymous):

He missed the 31 in 8x+y=31 when re-posting the question.

OpenStudy (anonymous):

8x+y= ? you missing

OpenStudy (anonymous):

no it was right ty

OpenStudy (anonymous):

31 nancy

OpenStudy (anonymous):

8x+y=31 time (-2) you get -16x-2y =-62 -x+2y=-6 -16x-2y =-62 ADD --------------- -17x = - 68 x =4

OpenStudy (anonymous):

substitution x =4 find y -x+2y=-6 -4+2y=-6 2y=-6+4 y= -2/2 y =-1

OpenStudy (anonymous):

time does that mean multiply the -2?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

Ty Nancy your a life saver=)